bobby567:

Complete Item List Chain Net Price Single Amount of Net Discount EQ Rate EQ Rate Trade Discount Price Rug $1,400 18/12 (A) (B) (C) (D) A. _____________ B. _____________ C. _____________ D. _____________

BlankSpace:

A bit confusing

BlankSpace:

Perhaps take a photo

bobby567:

Complete Item List Chain Net Price Single Amount of Net Discount EQ Rate EQ Rate Trade Discount Price Rug $1,400 18/12 (A) (B) (C) (D) A. _____________ B. _____________ C. _____________ D. _____________ idk if that helps but

Hoodmemes:

ye its helps

BlankSpace:

If you take a screenshot and upload it, that would be perfect if you can do that

bobby567:

idk how

Hoodmemes:

item:Rug List:1,400 Chain Discount:18/12 Net Price EQ Rate:(A)=unsolved Single Rate:(B)=unsolved Amount of trade discount:(C)=unsolved Net Price:(D)=unsolved A. _____________ B. _____________ C. _____________ D. _____________

Hoodmemes:

im thinkin this was what u was tryna type if im correct?

bobby567:

sure

Hoodmemes:

ah ok ima help

bobby567:

ok thank you

Hoodmemes:

Ight so 1 - 18% = 1 - 0.18 = 0.82,1 - 12% = 1 - 0.12 = 0.88,Amount trade discount (C)=0.82 *0.88 = 0.7216,(C) = 1 - 0.7216 = 100% - 72.16% = (blank%),Net price,(D) = $1400 * 0.7216 = ($bank)